OpenStudy (anonymous):

if g(x) =x^2+3x-1, find g(x+\[\Delta\]x)-g(x)/\[\Delta\]x

OpenStudy (anonymous):

hahaha definition of derivative

OpenStudy (anonymous):

umm whatt?

OpenStudy (anonymous):

are you in calculus?

OpenStudy (anonymous):

yes......

OpenStudy (anonymous):

that is the definition of a derivative of a function the answer will be 2x+3

OpenStudy (anonymous):

you will find a much more simpler way to solve this but uuntil then you must use this

jimthompson5910 (jim_thompson5910):

not quite, it will involve terms of delta as well

OpenStudy (anonymous):

but wouldnt there have to be a delta x in the middle of that?

OpenStudy (anonymous):

no the deltas will cancel i believe if i remember right

jimthompson5910 (jim_thompson5910):

you're thinking of evaluating the limit as delta --> 0

OpenStudy (anonymous):

well why the hell would they have you do this in calculus without the limit to 0?

jimthompson5910 (jim_thompson5910):

it's just the difference quotient, not the actual derivative it's a lead up to derivatives

OpenStudy (anonymous):

so there is a delta x in it? can one of u show how u guys r getting two different answers?

OpenStudy (anonymous):

o wow i never did it without it jim_thompson is right. You will have deltas however the whole point of this is to use it in this formula \[\lim_{\Delta x \rightarrow 0}\frac{ f(c + \Delta x)-f(c)}{\Delta x}\]

OpenStudy (anonymous):

ohh ok and i ended up with 2x+deltax+3 and thanksss

OpenStudy (zarkon):

not always...sometimes you are interested in the average rate of change and not the instantaneous rate of change

jimthompson5910 (jim_thompson5910):

\[\large g(x) = x^2+3x-1\] \[\large g(x+\Delta) = (x+\Delta)^2+3(x+\Delta)-1\] \[\large g(x+\Delta) = x^2+2x\Delta+\Delta^2+3x+3\Delta-1\] ----------------------------------------------------------------- \[\large \frac{g(x+\Delta)-g(x)}{\Delta}\] \[\large \frac{(x^2+2x\Delta+\Delta^2+3x+3\Delta-1)-(x^2+3x-1)}{\Delta}\] \[\large \frac{x^2+2x\Delta+\Delta^2+3x+3\Delta-1-x^2-3x+1}{\Delta}\] \[\large \frac{2x\Delta+\Delta^2+3\Delta}{\Delta}\] \[\large \frac{\Delta(2x+\Delta+3)}{\Delta}\] \[\large \frac{\cancel{\Delta}(2x+\Delta+3)}{\cancel{\Delta}}\] \[\large 2x+\Delta+3\] \[\large 2x+3+\Delta\] So \[\large \frac{g(x+\Delta)-g(x)}{\Delta}=2x+3+\Delta\]

OpenStudy (anonymous):

^